OpenStudy (anonymous):

solve for x. 3=log8+3logx

OpenStudy (luigi0210):

Condense the log on the right side

OpenStudy (luigi0210):

@TURITW don't do their work for them

OpenStudy (anonymous):

3=log8+3logx log8= log (2^3)=3log2 3=3log2+3logx 1=log2+logx

OpenStudy (anonymous):

use log a +log b=log ab